# EgoBlind: Towards Egocentric Visual Assistance for the Blind

### v2

@@ 5 Conclusion
We present the construction of a VideoQA dataset EgoBlind to reflect on the challenges towards egocentric visual assistance for the blind people. EgoBlind is the first of its kind in that the videos are token by the visually-impaired individuals from a first-person perspective, with the questions posed and verified by blind users to ensure close alignment with their true assistant needs. We further comprehensively benchmark the challenge with multiple prominent multimodal LLMs and unveil their significant limitations in various aspects. We conduct thorough analysis and share many key insights for future advancements in this direction. By formulating and bringing the challenge to the vision-language community, our primary goal is to push the MLLM research towards live egocentric visual assistance for the overwhelming number of visually impaired people around the world.
@@ A.2 Blind User Study
The cohort study (Figure 6(a)) exhibits substantial demographic heterogeneity, encompassing participants ranging from teenagers to older adults (Figure 6(b)). Male participants account for a dominant proportion of 73.3%, with young adults aged 18-30 constituting the largest age cohort. Notably, for most respondents, the duration of visual impairment corresponded with their biological age, suggesting congenital origin - a finding consistent with the observation that over 80% presented Grade 1 Blindness (total visual acuity $\leq$0.02 or visual field <5absent5<5< 5°). The remaining participants acquire visual impairment postnatally, though all reported living with the condition for long time.
@@
In outdoor navigation scenarios, such as finding destinations, crossing roads, or walking, blind individuals prioritize safety above all else (as shown in Figure 9). For instance, the question “Is it safe to cross the road?" receives the highest score of 5 from all respondents. Similarly, over 80% of respondents rate questions related to safety during movement, such as “How can I safely get through this entrance?" and “Is it safe to go straight?", with scores of 4 or higher. Blind individuals also place significant emphasis on directional and positional information. For questions about accessing public facilities, such as “How to reach the nearest sidewalk, zebra crossing, or blind road?", over 70% of respondents gave scores of 4 or higher. Direction-related questions like “Which direction should I take to reach the intersection?" and “Is the supermarket on my left?" received scores above 3 from more than 60% of respondents. However, respondents show less concern for questions about the characteristics of the surrounding environment, such as “Please describe the environment around me" and “Is traffic busy in this area?". Similarly, questions about weather conditions (“Is it sunny now?") or information about nearby pedestrians (“Who is the person standing in front?", “Are there many people around us gathered?", “Is there a pedestrian coming towards me?") received scores below 3 from more than 50% of respondents, indicating a general lack of interest. Interestingly, questions about obstacles did not receive as high a score as might be intuitively expected. For example, fewer than 50% of respondents positively rated questions like “Can you describe the obstacles ahead?" and “To avoid obstacles, which direction should I go?", while a large proportion gave neutral scores (3). This may be due to differences in the proficiency of using white canes among blind individuals, which influences their reliance on such information. Similarly, questions such as “Am I walking on a blind road” also exhibited a polarized response, with respondents either assigning very low scores (0 or 1) or very high scores (4 or 5). This polarization suggests that reliance on certain types of information may vary significantly depending on individual preferences and skills.
@@ A.3 Analysis of Videos
We analyze more details about the videos in EgoBlind. First, unlike the famous Ego4D [32] videos which are framed in a top-down view of sighted people via head-mounted camera devices (e.g., smart glasses), the videos in EgoBlind are captured by GoPros mounted in front of the blind users’ chests or mobile phones held in their hands. Specific challenges resulting from EgoBlind videos are: 1) The objects of interest are often off-center and out-of-focus due to blindness, as shown in Figure 12(a), 2) the objects are shown in challenging viewpoint (blind users tend to walk alongside a street curb or a building for safety.), and 3) The scene could be largely occluded by their canes, mobile phones, guide dogs or other people.
@@ B.1 Evaluation
Figure 13: Consistency between human and machine scoring.
@@ B.2 Benchmark Analysis
Figure 14 shows other common failure cases of the current Video-LLMs towards egocentric visual assistance for the blind. The example in Figure 14(a) demonstrates that the models are hard to perform effective live QA with streaming visual inputs. They often rely on the wrong frames for responses, even though we have explicitly prompted that the question is posed at the last frame (refer to our prompts at Table 9). The example in Figure 14(b) indicates that the models are poor at scene text recognition, as all model fail to answer the 3rd floor which was pressed by the user. The example also shows that the models are weak in performing temporal context reasoning. Because human can easily know the answer as the user moves her hand upward from button "1" to the 3rd button, even though the button "3" is partly blocked by the hand. Finally, the example in Figure 14(c) suggests that the models are struggling in linking the referential words (e.g.,“this") with the visual contents. This also necessitates spatio-temporal context understanding, e.g., following the hand motion and the object it interacted with.
@@ B.3 Other Investigations
We analyze more details about our investigations on single frame input, blind-specific prompt, and instruction tuning with EgoBlind training set.
B.3.1 Single Frame Input
Figure 15 shows that the overall model performances decline when multi-frame input is replaced with a single frame at the question timestamp. However, the degradations are relatively small (e.g., less than 1%). A breakdown by question type reveals that model accuracy drops significantly for questions related to “Tool Use” and “Navigation,” while it actually improves for “Information Reading” questions which account for nearly half of the dataset. As a result, the overall performance drop is mitigated by these gains. Figure 17(a) shows two cases where models fail to answer using only the question frame but succeed when provided with multiple frames. It reveals that multiple frames are helpful when the question-frame is not in good condition for answering, even if the question itself does not inherently require temporal understanding.
(a) InternVL2.5.
(b) Gemini 1.5 Flash.
(c) GPT-4o.
Figure 15: Uniform video sampling vs. single frame (Q-Frame) input at the question timestamp. The overall QA performance (average accuracy) declines slightly when replacing video with a single question frame. Specifically, for questions of “Navigation”, “Tool Use”, and “Communication”, video-level modeling is key for superior performance, while a single frame can better answer question related to information reading.
B.3.2 Blind-Aware Prompt
Figure 16 demonstrates that a blind-aware prompt are key for better performances. With a normal VQA prompt, the performances of all models decline remarkably for answering almost all types of questions. By analyzing the examples in Figure 17(b), we find that the models are either hard to understand the questions, or tend to give objectively correct answers about the visual contents but are helpfulless for the blind users.
(a) InternVL2.5.
(b) Gemini 1.5 Flash.
(c) GPT-4o.
Figure 16: Blind aware prompting vs. Normal VQA prompting. The model performances decline remarkably in answering almost all types of questions when substituting the blind-aware prompt (Table 9) with a normal vqa prompt (Table 8).
Figure 17: Visualization of predictions.
B.3.3 Instruction Tuning
@@
Implementation Details. The fine-tuning procedures are conducted on two NVIDIA A800 GPUs. Most hyper-parameters are kept at their default configurations as specified in the fine-tuning scripts. For InternVL2.5-8B, optimal performance is achieved by increasing the training epochs to 2 while finetuning only the MLP layer. Regarding LLaVA-OV, we employ LoRA with reduced rank dimensionality (r=16) to optimize computational efficiency. Due to constraints of compute resource, we truncate the maximum sampled frames to 16, and cape the model’s maximum sequence length to 8192, and freeze the vision encoder. Other refinements involve extending the number of epochs to 2, increasing gradient accumulation steps to 4, and adjusting weight decay to 0.05. Finally, for Qwen2.5-VL, we find that the best results are achieved by simply training the model 1 epoch while preserving other parameters unchanged.
Analysis. Figure 18 shows that the overall QA performances of the open-source models can be significantly improved by finetuning with EgoBlind training data. However, we find that the improvements are not consistent across different question types. For “Tool Use” questions, finetuning leads to a performance decline across all 3 models. We speculate that “Tool Use” relies more on procedural and instructional knowledge — areas where large language models (LLMs) excel — but requires less visual understanding. Fine-tuning may overwrite this prior knowledge, causing the models to overfit to the limited training data.
(a) Qwen2.5-VL.
(b) LLaVA-OV.
(c) InternVL-2.5-8B.
Figure 18: Instruction tuning with EgoBlind training data can significantly boost the overall QA performances but is detrimental for answering tool-use questions.
@@ Appendix D Societal Impacts
I want you to act as a blind person. I will provide you with several pictures each time and the pictures have two characteristics: 1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ‘communication and interaction’. Below I will give a detailed explanation of ‘communication and interaction’. Communication and Interaction: Describe the interaction between blind people and surrounding people/companions, and observe other people’s status information for multi-person collaborative activities. For example: Did my companion help me carry my luggage onto the conveyor belt? /Has my guide dog entered the elevator safely? /Who is talking to me? Here’s how you can accomplish the task: 1. use your imagination and only ask about ‘communication and interaction’ questions; 2.Please ask questions about each picture; 3. The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ‘I don’t know’; 6. Ask as many open-ended questions as possible, not just ’yes-no’ questions; 7. Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer. For example:’What is my companion doing? 8. ‘communication and interaction’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). 9. Please don’t ask bad examples like ‘What is my companion saying to me ?’/‘Is there an announcement about the next station or any other important information?’ Please generate the response with ’timestamp’, ’question’, ’answer’ and ’questoin_type’. The content of the key ’timestamp’ the time value corresponding to the current picture, for example: the first image corresponds to the first element in timestamp list, and so on. The timestamp value can be found here: ’f’ Timestamp list: time.

### v3

@@ 5 Conclusion
We present the construction of a VideoQA dataset EgoBlind to reflect on the challenges towards egocentric visual assistance for the blind. EgoBlind is the first of its kind in that the videos are taken by the visually-impaired individuals from a first-person perspective, with questions posed and verified by blind users to ensure close alignment with their true assistant needs. We further comprehensively benchmark the challenge with multiple prominent multimodal LLMs and unveil their significant limitations in various aspects. We conduct thorough analysis and share many key insights for future advancements in this direction. By formulating and bringing the challenge to the vision-language community, our primary goal is to push the MLLM research towards live egocentric visual assistance for the overwhelming number of visually impaired people around the world.
@@ A.2 Blind User Study
The cohort study (Figure 9(a)) exhibits substantial demographic heterogeneity, encompassing participants ranging from teenagers to older adults (Figure 9(b)). Male participants account for a dominant proportion of 73.3%, with young adults aged 18-30 constituting the largest age cohort. Notably, for most respondents, the duration of visual impairment corresponded with their biological age, suggesting congenital origin - a finding consistent with the observation that over 80% presented Grade 1 Blindness (total visual acuity $\leq$0.02 or visual field <5<5°). The remaining participants acquire visual impairment postnatally, though all reported living with the condition for long time.
@@
In outdoor navigation scenarios, such as finding destinations, crossing roads, or walking, blind individuals prioritize safety above all else (as shown in Figure 12). For instance, the question “Is it safe to cross the road?" receives the highest score of 5 from all respondents. Similarly, over 80% of respondents rate questions related to safety during movement, such as “How can I safely get through this entrance?" and “Is it safe to go straight?", with scores of 4 or higher. Blind individuals also place significant emphasis on directional and positional information. For questions about accessing public facilities, such as “How to reach the nearest sidewalk, zebra crossing, or blind road?", over 70% of respondents gave scores of 4 or higher. Direction-related questions like “Which direction should I take to reach the intersection?" and “Is the supermarket on my left?" received scores above 3 from more than 60% of respondents.
However, respondents show less concern for questions about the characteristics of the surrounding environment, such as “Please describe the environment around me" and “Is traffic busy in this area?". Similarly, questions about weather conditions (“Is it sunny now?") or information about nearby pedestrians (“Who is the person standing in front?", “Are there many people around us gathered?", “Is there a pedestrian coming towards me?") received scores below 3 from more than 50% of respondents, indicating a general lack of interest. Interestingly, questions about obstacles did not receive as high a score as might be intuitively expected. For example, fewer than 50% of respondents positively rated questions like “Can you describe the obstacles ahead?" and “To avoid obstacles, which direction should I go?", while a large proportion gave neutral scores (3). This may be due to differences in the proficiency of using white canes among blind individuals, which influences their reliance on such information. Similarly, questions such as “Am I walking on a blind road” also exhibited a polarized response, with respondents either assigning very low scores (0 or 1) or very high scores (4 or 5). This polarization suggests that reliance on certain types of information may vary significantly depending on individual preferences and skills.
@@ A.3 Analysis of Videos
We analyze more details about the videos in EgoBlind. First, unlike the famous Ego4D [35] videos which are framed in a top-down view of sighted people via head-mounted camera devices (e.g., smart glasses), the videos in EgoBlind are captured by GoPros mounted in front of the blind users’ chests
Figure 15: Consistency between human and machine scoring.
or mobile phones held in their hands. Specific challenges resulting from EgoBlind videos are: 1) The objects of interest are often off-center and out-of-focus due to blindness, as shown in Figure 16(a), 2) the objects are shown in challenging viewpoint (blind users tend to walk alongside a street curb or a building for safety.), and 3) The scene could be largely occluded by their canes, mobile phones, guide dogs or other people.
@@ B.2 Benchmark Analysis
Other Common Challenges. Figure 17 shows other common failure cases of the current Video-LLMs towards egocentric visual assistance for the blind. The example in Figure 17(a) demonstrates that the models are hard to perform effective live QA with streaming visual inputs. They often rely on the wrong frames for responses, even though we have explicitly prompted that the question is posed at the last frame (refer to our prompts at Table 12). The example in Figure 17(b) indicates that the models are poor at scene text recognition, as all model fail to answer the 3rd floor which was pressed by the user. The example also shows that the models are weak in performing temporal context reasoning. Because human can easily know the answer as the user moves her hand upward from button "1" to the 3rd button, even though the button "3" is partly blocked by the hand. Finally, the example in Figure 17(c) suggests that the models are struggling in linking the referential words (e.g.,“this") with the visual contents. This also necessitates spatio-temporal context understanding, e.g., following the hand motion and the object it interacted with.
@@ B.3 Other Investigations
Table 8: Overall QA accuracy (%) of user-agonistic data split vs. user-aware data split.
Models User-Agnostic User-Aware
Zero-shot Finetune Zero-shot Finetune
Qwen2.5-VL 45.5 50.2 $\uparrow$ 4.7 47.7 52.9 $\uparrow$ 5.2
LLaVA-OV 54.5 57.4 $\uparrow$ 2.9 56.4 62.0 $\uparrow$ 5.6
InternVL2.5-8B 53.5 58.1 $\uparrow$ 4.6 56.2 62.3 $\uparrow$ 6.1
Table 9: Results of category-specific prompting for each question group (half of test data).
Method category prompt Tool Info. Nav. Safe Com. Res. Overall
LLaVA-OV 58.4 54.1 37.2 63.8 59.4 54.0 54.2
50.6 53.2 30.2 64.1 56.2 52.9 52.2 $\downarrow$ 2.0
InternVL2.5-26B 74.0 56.0 47.7 51.9 46.9 56.3 54.6
66.2 55.5 52.8 43.0 53.1 63.2 53.1 $\downarrow$ 1.5
GPT-4o 61.0 59.6 54.3 60.3 46.9 69.0 59.4
79.2 63.3 58.8 58.6 50.0 64.4 62.2 $\uparrow$ 2.8
Table 10: Investigation of Chinese-specific elements (half of test data). Subtitles and scene texts (SText) play little role in EgoBlind, while all model performance decline significantly when translating the questions into Chinese (CHN).
Method Subt. SText CHN Tool Info. Nav. Safe Com. Res. Overall
Qwen2.5-VL 45.4 49.0 32.2 46.5 40.6 35.6 44.5
44.2 46.3 27.6 51.0 40.6 31.0 43.2 $\downarrow$ 1.3
42.9 48.5 33.7 45.5 46.9 43.7 44.8 $\uparrow$ 0.3
44.2 46.4 31.2 40.1 31.2 39.1 41.5 $\downarrow$ 3.0
LLaVA-OV 58.4 54.1 37.2 63.8 59.4 54.0 54.2
52.0 54.4 34.2 64.4 59.4 55.2 53.7 $\downarrow$ 0.5
52.0 56.4 35.2 62.5 53.1 54.0 54.1 $\downarrow$ 0.1
52.0 53.0 36.7 60.3 40.6 46.0 51.4 $\downarrow$ 2.8
InternVL2.5-26B 74.0 56.0 47.7 51.9 46.9 56.3 54.6
67.5 52.5 51.3 53.8 50.0 57.5 53.8 $\downarrow$ 0.8
62.3 57.4 48.7 49.7 53.1 55.2 54.2 $\downarrow$ 0.4
59.7 56.0 48.7 50.3 50.0 49.4 53.1 $\downarrow$ 1.5
GPT-4o 61.0 59.6 54.3 60.3 46.9 69.0 59.4
68.8 56.9 53.8 55.8 53.1 70.1 57.6 $\downarrow$ 1.8
63.6 59.0 50.8 53.2 56.2 62.1 56.7 $\downarrow$ 2.7
64.9 55.1 51.8 56.4 56.2 60.9 55.9 $\downarrow$ 3.5
@@
Implementation Details for Finetuning. The fine-tuning procedures are conducted on two NVIDIA A800 GPUs. Most hyper-parameters are kept at their default configurations as specified in the fine-tuning scripts. For InternVL2.5-8B, optimal performance is achieved by increasing the training epochs to 2 while finetuning only the MLP layer. Regarding LLaVA-OV, we employ LoRA with reduced rank dimensionality (r=16) to optimize computational efficiency. Due to constraints of compute resource, we truncate the maximum sampled frames to 16, and cape the model’s maximum sequence length to 8192, and freeze the vision encoder. Other refinements involve extending the number of epochs to 2, increasing gradient accumulation steps to 4, and adjusting weight decay to 0.05. Finally, for Qwen2.5-VL, we find that the best results are achieved by simply training the model 1 epoch while preserving other parameters unchanged.
Category-aware Prompting. We investigate whether designing category-specific prompt for each question group can enhance QA performance. To reduce API costs for closed-source models, the experiments are conducted on a randomly selected half of the test set. Results in Table 9 show that category-specific prompting (see Table 13) effectively improves GPT-4o’s performance, but not for other open-source models. We speculate that the open-source models are not strong enough to understand longer and complex prompts.
User-aware Dataset Split. Additionally, we re-split our EgoBlind dataset into train and test sets by ensuring videos token by the same user not appearing in both, i.e., half users in train and half users in test. We obtain both the zero-shot and finetuned results of three representative open-source models on the new test set in Table 8. The results show that finetuning can remarkably improve the model performance under different data splits. To our surprise, all models perform better with the user-aware split, regardless of zero-shot or finetuning. We speculate that some user videos and questions are not that challenging to answer.
Chinese Contents. While EgoBlind aims for “in-the-wild” visual assistance for the blind, the videos are majorly sourced from Chinese video sharing platforms. Hence, we study if the model performances are sensitive to Chinese-specific elements in the videos. Specifically, we first remove all subtitles from videos by using off-the-shelf subtitle remover  [^2] https://github.com/YaoFANGUK/video-subtitle-remover. Results in Table 10 show that all model performances drop slightly by 0.5% to 1.8%. A manual check of the new failure cases reveals that the vast majority (95.5% to 98.4% for different models) are irrelevant to subtitles and cloud be due to visual input variation resulted from subtitle removal (e.g., visual blur), not because of no subtitles. Alternatively, we extract scene texts from videos by using advanced OCR tool [^3] https://github.com/PaddlePaddle/PaddleOCR. The scene texts are used as additional prompting contents for MLLMs to make decisions. The results in Table 10 show that the additional scene texts do not help model predictions but slightly hurt the performance. The results and analyses suggest that Chinese subtitles and scene texts matter little in answering EgoBlind questions. Finally, we translate all questions into Chinese and find that all model performance decline significantly, reflecting that all models are weaker in understanding Chinese compared to English.
@@ Appendix D Societal Impacts
Table 13: Category-specific prompts. They are properly inserted into the QA prompts in Table 12.
Question Category Prompts
Tool Use The questions focus on tool use, helping the blind understand how to operate tools around them.
Information Reading The questions focus on information reading, helping the blind understand visible facts such as what their surroundings look like.
Navigation The questions focus on navigation, guiding the blind to move safely or find directions in their environment.
Safety Warnings The questions focus on safety warnings, reminding the blind of possible dangers of their surroundings.
Social Communication The questions focus on social communication, recognizing or describing people interacting with the blind.
Resource The questions focus on other resources, identifying people or things nearby that may offer help.
@@
I want you to act as a blind person. I will provide you with several pictures each time and the pictures have two characteristics: 1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about ‘communication and interaction’. Below I will give a detailed explanation of ‘communication and interaction’. Communication and Interaction: Describe the interaction between blind people and surrounding people/companions, and observe other people’s status information for multi-person collaborative activities. For example: Did my companion help me carry my luggage onto the conveyor belt? /Has my guide dog entered the elevator safely? /Who is talking to me? Here’s how you can accomplish the task: 1. use your imagination and only ask about ‘communication and interaction’ questions; 2.Please ask questions about each picture; 3. The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ‘I don’t know’; 6. Ask as many open-ended questions as possible, not just ’yes-no’ questions; 7. Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer. For example:’What is my companion doing? 8. ‘communication and interaction’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). 9. Please don’t ask bad examples like ‘What is my companion saying to me ?’/‘Is there an announcement about the next station or any other important information?’ Please generate the response with ’timestamp’, ’question’, ’answer’ and ’questoin_type’. The content of the key ’timestamp’ the time value corresponding to the current picture, for example: the first image corresponds to the first element in timestamp list, and so on. The timestamp value can be found here: ’f’ Timestamp list: time.